# Isometric immersions into three-dimensional unimodular metric Lie groups

### v1

@@ 3.4. Recovering the immersion from the angle functions
Let $\mathbb{R}^{3}$ endowed with the Euclidean metric, which satisfies $\mu_{1}=\mu_{2}=\mu_{3}=0$ and $\zeta=0$, so that (3.39) holds. We will see that all surfaces have angular companions even in the minimal case (in which $\psi\equiv 0$ and Theorem 3.8 does not apply). Let $\phi:\Sigma\to\mathbb{R}^{3}$ be any immersion and let $N$ be the unit normal. The antipodal map $\tau(x,y,z)=(-x,-y,-z)$ is an isometry, whence $\tau\circ\phi:\Sigma\to\mathbb{R}^{3}$ is also an isometric immersion with the same mean curvature but opposite angles when $-N=\tau_{*}N$ is chosen as unit normal. However, $\tau$ reverses the orientation, so the proper choice of normal for $\tau\circ\phi$ compatible with the orientation of $\Sigma$ is the opposite (i.e., $N=-\tau_{*}N$), which gives the same angle functions but opposite mean curvature as $\phi$ (note also that this antipodal immersion $\tau\circ\phi$ changes the signs of $T_{1},T_{2},T_{3}$). We deduce that $\phi$ and $\tau\circ\phi$ are angular companions.
@@
The case $\psi\equiv 0$ in $\mathbb{R}^{3}$ can be elaborated further. A minimal isometric immersion $\phi:\Sigma\to\mathbb{R}^{3}$ has Weierstrass data $(\mathcal{G},\eta)$, where $\mathcal{G}=\varphi(\nu_{1},\nu_{2},\nu_{3}):U\to\mathbb{C}$ is the stereographic projection of the Gauss map and $\eta$ is the height differential. If another minimal isometric immersion $\widetilde{\phi}:\Sigma\to\mathbb{R}^{3}$ has the same Gauss map, then its data are $(\mathcal{G},\widetilde{\eta})$ with possibly a different $\widetilde{\eta}$. In an isothermal coodinate $z$, we can express $\eta=f(z)\mathrm{d}z$ and $\widetilde{\eta}=\widetilde{f}(z)\mathrm{d}z$, and the intrinsic curvature of $\Sigma$ is well known to be $K=-16|\mathcal{G}^{\prime}|^{2}/(|f|^{2}(1+|\mathcal{G}|^{2})^{2})=-16|\mathcal{G}^{\prime}|^{2}/(|\widetilde{f}|^{2}(1+|\mathcal{G}|^{2})^{2})$, from where we deduce that $|\widetilde{f}|=|f|$ at non-flat points. Since $f$ and $\widetilde{f}$ are holomorphic, there is $\theta\in\mathbb{R}$ such that $\widetilde{f}=e^{i\theta}f$, and hence $\phi$ and $\widetilde{\phi}$ are in the same Bonnet family. Although this argument is local, it becomes global by the analiticity of minimal surfaces. Notice that, if all points happen to be flat, then $\phi$ and $\widetilde{\phi}$ are planes (with constant angle functions).
@@
where the last equality follows by expanding the gradients by means of the compatibility equations ((((v)))) and using the identity $JS+SJ=2HJ$. Equation (3.42) does not seem to have an easy resolution, even for cmc surfaces.
However, if we restrict again to vertical cylinders ($\nu_{3}=0$) and we assume $\kappa\neq 0$ (i.e., the space is not $\mathrm{Nil}_{3}$), we get that $\zeta=\frac{\kappa}{2\tau}\neq 0$, and hence (3.42) reduces to $\nabla H=0$. Therefore, the surface projects to a curve of constant curvature $2H$ in $\mathbb{M}^{2}(\kappa)$. If $04H^{2}+\kappa>$ 0, this can be seen as an Euclidean cylinder
@@
in the Cartan model. Again, $u$ and $v$ are orthogonal coordinates with $v$ running in the direction of $E_{3}$, and the induced metric reads $\frac{16r^{2}}{(4+\kappa r^{2})^{2}}(\mathrm{d}u^{2}+\mathrm{d}v^{2})$. The radius $r$ satisfies $04+\kappa r^{2}>$ 0 (so the surface actually lies in the Cartan model) and the mean curvature is given by $H=\frac{-4+\kappa r^{2}}{8r}$. The angular companion of this cylinder is nothing but a reparametrization of the same cylinder given by
@@
and combining (3.43) and (3.44) we conclude that $\widetilde{S}-J={\operatorname{Rot}_{\theta}}\circ{(S-J)}$. Taking into account that $\operatorname{Rot}_{\theta}=\frac{1-H^{2}}{1+H^{2}}\mathrm{id}-\frac{2H}{1+H^{2}}J$, this last relation can be rewritten as $\widetilde{S}-\widetilde{H}\,\mathrm{id}={\operatorname{Rot}_{\theta}}\circ{(S-H\,\mathrm{id})}$, that is, the traceless shape operator rotates a constant angle. We deduce that the angular companion of a cmc surface in $\mathbb{S}^{3}$ belongs to Lawson’s associate family, see [16, Thm. 8]. Indeed, these angular companions are twin immersions in the sense of Daniel, see [6, Thm. 5.14]. Note that our argument in particular shows that twin immersions have the same angle functions (it is not so well known that Lawson correspondence when acting on surfaces in $\mathbb{S}^{3}$ and $\mathbb{R}^{3}$ preserves the angle functions up to ambient isometries, see [14]).
Note also that twin $H$-immersions in $\mathbb{E}(\kappa,\tau)$ are not angular companions in general, because all cmc surfaces have twin immersions but not all of them satisfy (3.42). For instance, the parabolic helicoids $P_{\kappa,\tau,H}$ with $\kappa<0$, $\tau\neq 0$ and $04H^{2}+\kappa<$ 0 (as described in [12, Eqn. (2.7)]) have constant angle $\nu_{3}^{2}=\frac{4H^{2}+\kappa}{\kappa-4\tau^{2}}$, whence $\zeta=\frac{-2H^{2}}{\tau}$ is a nonzero constant. Since $T_{3}$ never vanishes in $P_{\kappa,\tau,H}$, we conclude that parabolic helicoids do not satisfy (3.42).
@@ 3.5. Constant angles and totally geodesic surfaces
00 0 $=-\epsilon_{1}\epsilon_{2}\epsilon_{3}\det\left(\begin{matrix}\epsilon_{1}\nu_{1}&\epsilon_{2}\nu_{2}&\epsilon_{3}\nu_{3}\\ c_{1}\nu_{1}&c_{2}\nu_{2}&c_{3}\nu_{3}\\ (\epsilon_{2}c_{2}-\epsilon_{3}c_{3})\nu_{2}\nu_{3}&(\epsilon_{3}c_{3}-\epsilon_{1}c_{1})\nu_{3}\nu_{1}&(\epsilon_{1}c_{1}-\epsilon_{2}c_{2})\nu_{1}\nu_{2}\end{matrix}\right)$
@@
Conversely, suppose that the immersion has constant angles $\nu_{1},\nu_{2},\nu_{3}\in\mathbb{R}$, so the normal reads $N=\sum_{\alpha=1}^{3}\epsilon_{\alpha}\nu_{\alpha}E_{\alpha}$. Assume first that $\nu_{3}\neq 0$, we can consider $Y_{1}=\nu_{3}E_{1}-\nu_{1}E_{3}$ and $Y_{2}=\nu_{3}E_{2}-\nu_{2}E_{3}$, which form a tangent frame in $\Sigma$ and satisfy $\langle[Y_{1},Y_{2}],N\rangle=\nu_{3}\zeta$. As $\nu_{3}\neq 0$, we deduce that the distribution $\Delta=\operatorname{span}\{Y_{1},Y_{2}\}$ (orthogonal to $N$) is integrable if and only if $\zeta=0$. If, on the contrary, we have $\nu_{3}=0$, the discussion is similar by defining $Y_{1}=\nu_{2}E_{1}-\nu_{1}E_{2}$ and $Y_{2}=E_{3}$ instead. Either way, since $N$ is a unit normal for the foliation consisting of the integral surfaces of $\Delta$ and $\operatorname{div}(N)=0$ (note that (2.4) implies that $\operatorname{div}(E_{\alpha})=0$ for all $\alpha\in\{1,2,3\}$), we conclude that $\Sigma$ is minimal. This argument also proves the last assertion in the statement. Note also that (ii) and (iii) are known to be equivalent in general (e.g., see [23, Lem. 3.9]). ∎
@@
We also remark that, in the Riemannian case, a classification of 2-dimensional Lie subgroups of $G$ can be found in [23]. In the case $\dim(\operatorname{Iso}(G))=3$, the connected groups of isometries of $G$ that act on $G$ with orbits of codimension one (namely, giving rise to extrinsically homogeneous surfaces) are nothing but the $2$-dimensional Lie subgroups of $G$. The classification of the corresponding cohomogeneity one actions up to orbit equivalence can be found in [8, Tab. 1], from where the classification of homogeneous surfaces up to congruence follows.
The family of surfaces with constant angles turns out to contain all totally geodesic surfaces in our unimodular metric Lie groups. This was proved by Tsukada [34] in the Riemannian case (see also [20, Prop. 3.4]). Here, we will provide a different proof of this result that also applies to the Lorentzian case.
@@
If the immersion is totally geodesic, Codazzi equation ((((ii)))) implies that the vector field $T$ given by (3.8) is identically zero, or equivalently
@@
The inner products of (3.48) and $T_{1},T_{2},T_{3}$ can be computed by means of the algebraic relations ((((iii)))) and yields the following equations:
@@
where $\varrho=(\epsilon_{1}\mu_{2}\mu_{3}\nu_{1}^{2}+\epsilon_{2}\mu_{3}\mu_{1}\nu_{2}^{2}+\epsilon_{3}\mu_{1}\mu_{2}\nu_{3}^{2})\epsilon_{3}$. If there are points where no angle function is zero, we get $\mu_{2}\mu_{3}=\mu_{3}\mu_{1}=\mu_{1}\mu_{2}=\varrho$, so the space has constant sectional curvature by Proposition 2.2, which is discarded in the statement. Moreover, if two of the angle functions vanish on (an open subset of) $\Sigma$, namely $\nu_{1}=\nu_{2}=0$ with no loss of generality, we get $\nu_{3}=\pm 1$, whence $\Sigma$ is an integral surface of the distribution $\operatorname{span}\{E_{1},E_{2}\}$, which is integrable if and only if $c_{3}=0$. From the compatibility equations ((((v)))), we get $\mu_{2}\nu_{3}T_{2}=\mu_{3}\nu_{2}T_{3}$ and $\mu_{1}\nu_{3}T_{1}=\mu_{3}\nu_{1}T_{3}$, and this implies $\mu_{1}T_{1}=\mu_{2}T_{2}=0$. As $T_{1}$ and $T_{2}$ do not vanish, it must be $\mu_{1}=\mu_{2}=0$, i.e., $\epsilon_{1}c_{1}=\epsilon_{2}c_{2}$ and the sectional curvature is constant again by Proposition 2.2.
We are left with the case just one of the angle functions vanish in (an open subset of) $\Sigma$, namely $\nu_{3}=0$ and $\nu_{1},\nu_{2}\neq 0$ with no loss of generality. Equation (3.49) tells us that $\varrho=\mu_{1}\mu_{3}=\mu_{3}\mu_{2}$, whence $\mu_{1}=\mu_{2}$ or $\mu_{3}=0$. Since $\nabla\nu_{3}=0$ and $S\equiv 0$, the compatibility equation ((((v)))) guarantees that $\mu_{2}\nu_{1}T_{2}=\mu_{1}\nu_{2}T_{1}$, so that $\eta=\mu_{2}\nu_{1}E_{2}-\mu_{1}\nu_{2}E_{1}$ is normal to the immersion. In particular, $\eta$ must be collinear with $N=\epsilon_{1}\nu_{1}E_{1}+\epsilon_{2}\nu_{2}E_{2}$, which amounts to $\epsilon_{1}\mu_{2}\nu_{1}^{2}+\epsilon_{2}\mu_{1}\nu_{2}^{2}=0$. This last equation along with $\epsilon_{1}\nu_{1}^{2}+\epsilon_{2}\nu_{2}^{2}=\epsilon_{3}$ form a linear system in the unknowns $\{\nu_{1}^{2},\nu_{2}^{2}\}$, which is compatible if and only if $\mu_{1}\neq\mu_{2}$, in which case has a unique solution. This rules out the case $\mu_{1}=\mu_{2}$, whence it must be $\mu_{3}=0$ (i.e., $\epsilon_{3}c_{3}=\epsilon_{1}c_{1}+\epsilon_{2}c_{2}$). The aforesaid unique solution gives constant angles
@@
Notice that the foliations by totally geodesic surfaces spanned by both distributions in Theorem 3.16 could be congruent to each other, see [8, Rmk. 4.1].
@@ 4. Bianchi-Cartan-Vrănceanu spaces
Observe that, in order to compute the $\mathfrak{so}^{\epsilon}_{3}$-symmetric matrix of $1$-forms $\Theta=\Omega+L(M)$ a priori we need the intrinsic forms $\omega^{i}_{j}$, with $i,j\in\{1,2\}$, the shape operator $S$ (encoded in $\omega^{1}_{3}$ and $\omega^{2}_{3}$), and all entries of the matrix $M$. If $\Sigma$ is simply connected, then the equation $M^{-1}\mathrm{d}M=\Theta$ has a solution $M$ if and only if $\Theta$ has zero Darboux derivative, i.e., $\mathrm{d}\Theta+\tfrac{1}{2}[\Theta,\Theta]=0$ (this is a classical result in the theory of moving frames, see [30, Ch. 3]). In this section, we will prescribe $S$ as a fundamental datum as Daniel does (in contrast with our previous results in Section 3), which is the same as prescribing $\Omega$, and additionally we would like to have $L(M)$ without knowing all the information contained in $M$. Next result shows that we are forced to assume $\dim(\operatorname{Iso}(G))\geq 4$.

### v2

@@ 3.4. Recovering the immersion from the angle functions
Let $\mathbb{R}^{3}$ endowed with the Euclidean metric, which satisfies $\mu_{1}=\mu_{2}=\mu_{3}=0$ and $\zeta=0$, so that (3.39) holds. We will see that all surfaces have angular companions even in the minimal case (in which $\psi\equiv 0$ and Theorem 3.8 does not apply). Let $\phi:\Sigma\to\mathbb{R}^{3}$ be any immersion and let $N$ be the unit normal. The antipodal map $\tau(x,y,z)=(-x,-y,-z)$ is an isometry, whence $\tau\circ\phi:\Sigma\to\mathbb{R}^{3}$ is also an isometric immersion with the same mean curvature but opposite angles when $N=\tau_{*}N$ is chosen as unit normal. However, $\tau$ reverses the orientation, so the proper choice of normal for $\tau\circ\phi$ compatible with the orientation of $\Sigma$ is the opposite (i.e., $N=-\tau_{*}N$), which gives the same angle functions but opposite mean curvature as $\phi$ (note also that this antipodal immersion $\tau\circ\phi$ changes the signs of $T_{1},T_{2},T_{3}$). We deduce that $\phi$ and $\tau\circ\phi$ are angular companions.
@@
[…]
@@
where the last equality follows by expanding the gradients by means of the compatibility equations ((v)) and using the identity $JS+SJ=2HJ$. Equation (3.42) does not seem to have an easy resolution, even for cmc surfaces.
However, if we restrict again to vertical cylinders ($\nu_{3}=0$) and we assume $\kappa\neq 0$ (i.e., the space is not $\mathrm{Nil}_{3}$), we get that $\zeta=\frac{\kappa}{2\tau}\neq 0$, and hence (3.42) reduces to $\nabla H=0$. Therefore, the surface projects to a curve of constant curvature $2H$ in $\mathbb{M}^{2}(\kappa)$. If $4H^{2}+\kappa>0$, this can be seen as an Euclidean cylinder
@@
in the Cartan model. Again, $u$ and $v$ are orthogonal coordinates with $v$ running in the direction of $E_{3}$, and the induced metric reads $\frac{16r^{2}}{(4+\kappa r^{2})^{2}}(\mathrm{d}u^{2}+\mathrm{d}v^{2})$. The radius $r$ satisfies $4+\kappa r^{2}>0$ (so the surface actually lies in the Cartan model) and the mean curvature is given by $H=\frac{-4+\kappa r^{2}}{8r}$. The angular companion of this cylinder is nothing but a reparametrization of the same cylinder given by
@@
and combining (3.43) and (3.44) we conclude that $\widetilde{S}-J={\operatorname{Rot}_{\theta}}\circ{(S-J)}$. Taking into account that $\operatorname{Rot}_{\theta}=\frac{1-H^{2}}{1+H^{2}}\mathrm{id}-\frac{2H}{1+H^{2}}J$, this last relation can be rewritten as $\widetilde{S}-\widetilde{H}\,\mathrm{id}={\operatorname{Rot}_{\theta}}\circ{(S-H\,\mathrm{id})}$, that is, the traceless shape operator rotates a constant angle. We deduce that the angular companion of a cmc surface in $\mathbb{S}^{3}$ belongs to Lawson’s associate family, see [17, Thm. 8]. Indeed, these angular companions are twin immersions in the sense of Daniel, see [7, Thm. 5.14]. Note that our argument in particular shows that twin immersions have the same angle functions (it is not so well known that Lawson correspondence when acting on surfaces in $\mathbb{S}^{3}$ and $\mathbb{R}^{3}$ preserves the angle functions up to ambient isometries, see [15]). It is also important to point out that twin $H$-surfaces in $\mathbb{S}^{3}$ are not related by ambient isometries (the above computations yield a nontrivial rotation of the traceless shape operators). Explicit examples of non globally congruent twin $H$-surfaces are the horizontal $H$-tubes, see [20, Thm. 5.5].
Note also that twin $H$-immersions in $\mathbb{E}(\kappa,\tau)$ are not angular companions in general, because all cmc surfaces have twin immersions but not all of them satisfy (3.42). For instance, the parabolic helicoids $P_{\kappa,\tau,H}$ with $\kappa<0$, $\tau\neq 0$ and $4H^{2}+\kappa<0$ (as described in [11, Eqn. (2.7)]) have constant angle $\nu_{3}^{2}=\frac{4H^{2}+\kappa}{\kappa-4\tau^{2}}$, whence $\zeta=\frac{-2H^{2}}{\tau}$ is a nonzero constant. Since $T_{3}$ never vanishes in $P_{\kappa,\tau,H}$, we conclude that parabolic helicoids do not satisfy (3.42).
@@ 3.5. Constant angles and totally geodesic surfaces
$0$ $=-\epsilon_{1}\epsilon_{2}\epsilon_{3}\det\left(\begin{matrix}\epsilon_{1}\nu_{1}&\epsilon_{2}\nu_{2}&\epsilon_{3}\nu_{3}\\ c_{1}\nu_{1}&c_{2}\nu_{2}&c_{3}\nu_{3}\\ (\epsilon_{2}c_{2}-\epsilon_{3}c_{3})\nu_{2}\nu_{3}&(\epsilon_{3}c_{3}-\epsilon_{1}c_{1})\nu_{3}\nu_{1}&(\epsilon_{1}c_{1}-\epsilon_{2}c_{2})\nu_{1}\nu_{2}\end{matrix}\right)$
@@
Conversely, suppose that the immersion has constant angles $\nu_{1},\nu_{2},\nu_{3}\in\mathbb{R}$, so the normal reads $N=\sum_{\alpha=1}^{3}\epsilon_{\alpha}\nu_{\alpha}E_{\alpha}$. Assume first that $\nu_{3}\neq 0$, we can consider $Y_{1}=\nu_{3}E_{1}-\nu_{1}E_{3}$ and $Y_{2}=\nu_{3}E_{2}-\nu_{2}E_{3}$, which form a tangent frame in $\Sigma$ and satisfy $\langle[Y_{1},Y_{2}],N\rangle=\nu_{3}\zeta$. As $\nu_{3}\neq 0$, we deduce that the distribution $\Delta=\operatorname{span}\{Y_{1},Y_{2}\}$ (orthogonal to $N$) is integrable if and only if $\zeta=0$. If, on the contrary, we have $\nu_{3}=0$, the discussion is similar by defining $Y_{1}=\nu_{2}E_{1}-\nu_{1}E_{2}$ and $Y_{2}=E_{3}$ instead. Either way, since $N$ is a unit normal for the foliation consisting of the integral surfaces of $\Delta$ and $\operatorname{div}(N)=0$ (note that (2.4) implies that $\operatorname{div}(E_{\alpha})=0$ for all $\alpha\in\{1,2,3\}$), we conclude that $\Sigma$ is minimal. This argument also proves the last assertion in the statement. Note also that (ii) and (iii) are known to be equivalent in general (e.g., see [24, Lem. 3.9]). ∎
@@
We also remark that, in the Riemannian case, a classification of 2-dimensional Lie subgroups of $G$ can be found in [24]. In the case $\dim(\operatorname{Iso}(G))=3$, the connected groups of isometries of $G$ that act on $G$ with orbits of codimension one (namely, giving rise to extrinsically homogeneous surfaces) are nothing but the $2$-dimensional Lie subgroups of $G$. The classification of the corresponding cohomogeneity one actions up to orbit equivalence can be found in [10, Tab. 1], from where the classification of homogeneous surfaces up to congruence follows.
The family of surfaces with constant angles turns out to contain all totally geodesic surfaces in our unimodular metric Lie groups. This was proved by Tsukada [35] in the Riemannian case (see also [21, Prop. 3.4]). Here, we will provide a different proof of this result that also applies to the Lorentzian case.
@@
If the immersion is totally geodesic, Codazzi equation ((ii)) implies that the vector field $T$ given by (3.8) is identically zero, or equivalently
@@
The inner products of (3.48) and $T_{1},T_{2},T_{3}$ can be computed by means of the algebraic relations ((iii)) and yields the following equations:
@@
where $\varrho=(\epsilon_{1}\mu_{2}\mu_{3}\nu_{1}^{2}+\epsilon_{2}\mu_{3}\mu_{1}\nu_{2}^{2}+\epsilon_{3}\mu_{1}\mu_{2}\nu_{3}^{2})\epsilon_{3}$. If there are points where no angle function is zero, we get $\mu_{2}\mu_{3}=\mu_{3}\mu_{1}=\mu_{1}\mu_{2}=\varrho$, so the space has constant sectional curvature by Proposition 2.2, which is discarded in the statement. Moreover, if two of the angle functions vanish on (an open subset of) $\Sigma$, namely $\nu_{1}=\nu_{2}=0$ with no loss of generality, we get $\nu_{3}=\pm 1$, whence $\Sigma$ is an integral surface of the distribution $\operatorname{span}\{E_{1},E_{2}\}$, which is integrable if and only if $c_{3}=0$. From the compatibility equations ((v)), we get $\mu_{2}\nu_{3}T_{2}=\mu_{3}\nu_{2}T_{3}$ and $\mu_{1}\nu_{3}T_{1}=\mu_{3}\nu_{1}T_{3}$, and this implies $\mu_{1}T_{1}=\mu_{2}T_{2}=0$. As $T_{1}$ and $T_{2}$ do not vanish, it must be $\mu_{1}=\mu_{2}=0$, i.e., $\epsilon_{1}c_{1}=\epsilon_{2}c_{2}$ and the sectional curvature is constant again by Proposition 2.2.
We are left with the case just one of the angle functions vanish in (an open subset of) $\Sigma$, namely $\nu_{3}=0$ and $\nu_{1},\nu_{2}\neq 0$ with no loss of generality. Equation (3.49) tells us that $\varrho=\mu_{1}\mu_{3}=\mu_{3}\mu_{2}$, whence $\mu_{1}=\mu_{2}$ or $\mu_{3}=0$. Since $\nabla\nu_{3}=0$ and $S\equiv 0$, the compatibility equation ((v)) guarantees that $\mu_{2}\nu_{1}T_{2}=\mu_{1}\nu_{2}T_{1}$, so that $\eta=\mu_{2}\nu_{1}E_{2}-\mu_{1}\nu_{2}E_{1}$ is normal to the immersion. In particular, $\eta$ must be collinear with $N=\epsilon_{1}\nu_{1}E_{1}+\epsilon_{2}\nu_{2}E_{2}$, which amounts to $\epsilon_{1}\mu_{2}\nu_{1}^{2}+\epsilon_{2}\mu_{1}\nu_{2}^{2}=0$. This last equation along with $\epsilon_{1}\nu_{1}^{2}+\epsilon_{2}\nu_{2}^{2}=\epsilon_{3}$ form a linear system in the unknowns $\{\nu_{1}^{2},\nu_{2}^{2}\}$, which is compatible if and only if $\mu_{1}\neq\mu_{2}$, in which case has a unique solution. This rules out the case $\mu_{1}=\mu_{2}$, whence it must be $\mu_{3}=0$ (i.e., $\epsilon_{3}c_{3}=\epsilon_{1}c_{1}+\epsilon_{2}c_{2}$). The aforesaid unique solution gives constant angles
@@
Notice that the foliations by totally geodesic surfaces spanned by both distributions in Theorem 3.16 could be congruent to each other, see [10, Rmk. 4.1].
@@ 4. Bianchi-Cartan-Vrănceanu spaces
Observe that, in order to compute the $\mathfrak{so}^{\epsilon}_{3}$-symmetric matrix of $1$-forms $\Theta=\Omega+L(M)$ a priori we need the intrinsic forms $\omega^{i}_{j}$, with $i,j\in\{1,2\}$, the shape operator $S$ (encoded in $\omega^{1}_{3}$ and $\omega^{2}_{3}$), and all entries of the matrix $M$. If $\Sigma$ is simply connected, then the equation $M^{-1}\mathrm{d}M=\Theta$ has a solution $M$ if and only if $\Theta$ has zero Darboux derivative, i.e., $\mathrm{d}\Theta+\tfrac{1}{2}[\Theta,\Theta]=0$ (this is a classical result in the theory of moving frames, see [31, Ch. 3]). In this section, we will prescribe $S$ as a fundamental datum as Daniel does (in contrast with our previous results in Section 3), which is the same as prescribing $\Omega$, and additionally we would like to have $L(M)$ without knowing all the information contained in $M$. Next result shows that we are forced to assume $\dim(\operatorname{Iso}(G))\geq 4$.